# Introducing a Physics-informed Deep Learning Framework for Bridge Scour Prediction

### v1

@@ Abstract
This paper introduces scour physics-informed neural network algorithms (SPINNs), a hybrid physics-data-driven framework for bridge scour prediction using deep learning developed based on historical scour monitoring data. SPINNs integrate physics-based empirical equations into neural networks as supplementary loss components. We examined three architectures: LSTM, CNN, and NLinear. While CNN and LSTM have shown competitive real-time scour forecasting in previous studies, NLinear with a simple architecture demonstrated the highest accuracy and significantly lower computational cost. Despite varying error reduction margins across different base models and bridges, SPINNs showed promising scour prediction and generally outperformed pure data-driven models. In some bridge cases, SPINN reduced forecasting errors by up to 50%. In this study, we also explored generalised models for bridge clusters by aggregating training datasets from multiple bridges in Alaska. Bridge/site-specific SPINNs incorporating HEC18 and time-dependent empirical equations provided more accurate predictions than SPINNs with generalised time-dependent equations. The three empirical equations derived from SPINN training in this study showed reasonable accuracy in estimating maximum scour depth. These deep learning derived empirical models can provide more accurate and reliable scour predictions than traditional HEC-18, particularly in scenarios lacking site-specific scour data. Comparing the HEC-18 model with both SPINNs and pure deep learning models highlights a substantial improvement in scour prediction accuracy, indicating a promising future for these hybrid machine learning methodologies for bridge scour design and maintenance.
@@ 2.3 Scour Physics-Informed Deep Learning Framework
We introduce Scour Physics-informed Neural Networks (SPINN) as a new physics-informed deep learning framework for scour prediction. This framework integrates an additional physical loss term into the objective function, guiding the deep learning models to learn from the specified empirical equation. The overall architecture of SPINN is depicted in Fig. 1.

### v2

@@ Abstract
This paper introduces scour physics-informed neural networks (SPINNs), a hybrid physics-data-driven framework for bridge scour prediction using deep learning. SPINNs are developed based on historical scour monitoring data and integrate physics-based empirical equations into neural networks as supplementary loss components. We incorporated three architectures: LSTM, CNN, and NLinear as the base data-driven model. Despite varying performance across different base models and bridges, SPINNs overall outperformed pure data-driven models. In some bridge cases, SPINN reduced forecasting errors by up to 50 percent. In this study, we also explored general models for bridge clusters, trained by aggregating datasets across multiple bridges in a region. The pure data-driven models mostly benefited from this approach, in particular bridges with limited data. However, bridge-specific SPINNs provided more accurate predictions than general SPINNs for almost all case studies. Also, the time-dependent empirical equations derived from SPINNs showed reasonable accuracy in estimating maximum scour depth, providing more accurate predictions compared to HEC-18. Comparing both SPINNs and pure deep learning models with traditional HEC-18 equation indicates substantial improvements in scour prediction accuracy. This study can pave the way for hybrid physics-machine learning methodologies to be implemented for bridge scour design and maintenance.
@@ 2.3 Scour Physics-Informed Deep Learning Framework
We introduce Scour Physics-informed Neural Network (SPINN) as a new physics-informed deep learning framework for scour prediction. This framework integrates an additional physical loss term into the objective function, guiding the deep learning models to learn from the specified empirical equation. The overall architecture of SPINN is depicted in Fig. 1.